OpenStudy (anonymous):

i know that Number of subsets of a set \(S\) of size n , is given by the binomial sum > **\(\sum_{k=1} ^n \binom{n}{k}=2^n\)** , n=1,2,3,.. elements how we could conclude of prove this formula ? how to prove that this gives the number of subsets . where did this formula come from ? **PS:-** i know how to prove \(\sum_{k=0} ^n \binom{n}{k}=2^n\)

OpenStudy (rational):

use induction

OpenStudy (anonymous):

no no no -.-

OpenStudy (rational):

and there are dozens of other proofs too, but i have worked this using induction recently... so...

OpenStudy (anonymous):

hmm what i need is how to get this http://prntscr.com/4j52yx

OpenStudy (rational):

Okay, lets do an informal proof using counting argument

OpenStudy (anonymous):

i know all proofs that shows \(\sum_{k=0} ^n \binom{n}{k}=2^n\) i wanna know where did it come from

OpenStudy (anonymous):

ok show me.

OpenStudy (rational):

you want a proof for number of subsets ? or the binomial identity : \(\sum_{k=0} ^n \binom{n}{k}=2^n\) ?

OpenStudy (kirbykirby):

Try using counting arguments. If you "expand" the summation as a sum of binomial coefficients, hopefully it will make sense to justify each of those combinations. Then you just "contract" the sum of these binomial coefficients using the binomial theorem to get 2^n

OpenStudy (anonymous):

i wanna prove number of subsets , like the number of subsets of a finite set.

OpenStudy (anonymous):

seems no one get my question -.- every one trying to teach me how binomial (1+1)^n goes to sum of (n,k) which i already know.

OpenStudy (aum):

If there are "n" elements in a set, you can form the following subsets: Select one item at a time. There will be nC1 subsets. Select two items at a time. There will be nC2 subsets. .... Select 'n' items at a time. There will be nCn subsets. Add them all.

OpenStudy (rational):

Notice that you can either include a particular element in your subset or not - So how many choices do you have for each element ?

OpenStudy (anonymous):

hmmm same idea aum

OpenStudy (anonymous):

ok so its about choises ? like how many sets i can get from n elements without repeating ?! OH my , thats exactly what i was looking for xD

OpenStudy (rational):

Thats not my question. Maybe lets work this with a general example. Suppose your master set \(\large M\) contains \(n\) elements

OpenStudy (rational):

Next, consider a subset \(\large A\)

OpenStudy (rational):

Any element in in \(\large M\) has exactly \(\large 2\) choices : 1) either the element exists in subset \(A\) 2) or it doesn't exist in subset \(A\)

OpenStudy (rational):

So total number of ways a subset can be formed = 2x2x...n times = 2^n

OpenStudy (anonymous):

i see mm

OpenStudy (anonymous):

that was quick

OpenStudy (rational):

its an one liner induction proof is more beautiful though

OpenStudy (anonymous):

ok show me

OpenStudy (kirbykirby):

I was going with the same idea as @aum which seemed fairly simple.

OpenStudy (rational):

will do induction some other time have u noticed that the previous counting argument is very much analogous to finding number of divisors of a number

OpenStudy (anonymous):

i know it its cool i know hmmm i was freaked out to see order stuff of subgroup -.-

OpenStudy (rational):

if \(\large n = p_1^{e_1} p_2^{e_2} \ldots p_r^{e_r} \) number of divisors = \((e_1+1)(e_2+1)\cdots (e_r+1)\)

OpenStudy (rational):

somehow its tricky to to see the exact connection, but you can see that both proofs are similar

OpenStudy (rational):

what about the order ?

OpenStudy (anonymous):

there is 500 page book i have to review this week lol im only lost how to start hehehe

OpenStudy (rational):

this link has induction proof http://www.artofproblemsolving.com/Wiki/index.php/Power_set

OpenStudy (anonymous):

ok cool , i know that too

OpenStudy (rational):

Oh so you did not open the textbook till now is it

OpenStudy (anonymous):

hmm haha clear , yes

OpenStudy (anonymous):

its like i know everything , but know nothing random knowllege controling my head but its like without direction somehow >.< idk if im making any sense

OpenStudy (anonymous):

not used to do it on my own.

OpenStudy (anonymous):

+ my teacher dont reply to my emails xD

OpenStudy (rational):

elderly advice : you should study from the start and stick to some timetable; if you study randomly, your grades also will be random :P

OpenStudy (anonymous):

thanks for the spirits up lol

OpenStudy (rational):

just saying lol one week=7days is good enough time to master anything, including wiles proof of fermat's last theorem ! good luck !!

OpenStudy (anonymous):

>.< u dnt know abstract algebra is fully

OpenStudy (rational):

you took abstract before right ? why are u studying again ?

OpenStudy (anonymous):

\[2^n=(1+1)^n\] Apply the binomial theorem.

OpenStudy (anonymous):

thank you all ^^